# Rethinking Patch Dependence for Masked Autoencoders

### v1

@@ Abstract
In this work, we re-examine inter-patch dependencies in the decoding mechanism of masked autoencoders (MAE). We decompose this decoding mechanism for masked patch reconstruction in MAE into self-attention and cross-attention. Our investigations suggest that self-attention between mask patches is not essential for learning good representations. To this end, we propose a novel pretraining framework: Cross-Attention Masked Autoencoders (CrossMAE). CrossMAE’s decoder leverages only cross-attention between masked and visible tokens, with no degradation in downstream performance. This design also enables decoding only a small subset of mask tokens, boosting efficiency. Furthermore, each decoder block can now leverage different encoder features, resulting in improved representation learning. CrossMAE matches MAE in performance with 2.5 to 3.7$\times$ less decoding compute. It also surpasses MAE on ImageNet classification and COCO instance segmentation under the same compute. Code and models: https://crossmae.github.io.
*Equal ${}^{\dagger}$qual advising.
@@ 1 Introduction
Masked image modeling [46, 31, 60, 5] has emerged as a pivotal unsupervised learning technique in computer vision. One such recent work following this paradigm is masked autoencoders (MAE): given only a small, random subset of visible image patches, the model is tasked to reconstruct the missing pixels. By operating strictly on this small subset of visible tokens, MAE can efficiently pre-train high-capacity models on large-scale vision datasets, demonstrating impressive results on a wide array of downstream tasks [34, 39, 49].
Figure 1: Method Overview. (A) Masked autoencoder (MAE) starts by masking random patches of the input image. (B) To reconstruct a mask token (i.e. marked by the blue star), MAE attends to both the masked tokens (B.Left) and the visible tokens (B.Right). A quantitative comparison over the ImageNet validation set shows that the masked tokens in MAE disproportionally attend to the visible tokens (1.42 vs 0.39), questioning the necessity of attention within mask tokens. (C) We propose CrossMAE, where each mask token can only attend to the visible tokens. (D) CrossMAE is equivalent to or better than MAE (Tabs. 1 and 3), with significantly less decoder FLOPS (2.5x lower).
Figure 2: Example reconstructions of ImageNet validation images. For each set of 5 images, from left to right, are the original image, masked image with a mask ratio of 75%, MAE [31], CrossMAE (trained to reconstruct 25% of image tokens, or 1/3 of the mask tokens), and CrossMAE (trained to reconstruct all masked tokens). Since CrossMAE does not reconstruct them, all model outputs have the visible patches overlaid. Intriguingly, CrossMAE, when trained for partial reconstruction, can decode all mask tokens in one forward pass (shown above), which deviates from its training methodology. Its comparable reconstruction quality to full-image-trained models suggests that full-image reconstruction might not be essential for effective representation learning.
@@
The MAE framework uses multi-headed self-attention throughout the model to perform the self-supervised reconstruction task, where the masked and visible tokens not only attend to each other but also to themselves, to generate a holistic and contextually aware representation. Yet, the mask tokens themselves do not contain information. Intuitively, allowing the exchange of information within nearby mask tokens is supposed to allow the model to synthesize a more consistent image; however, is dense self-attention necessary for learning good representation? We analyze the two parallel components involved in decoding each mask token: self-attention with other mask tokens, as well as cross-attention to the encoded visible tokens. If MAE relies on the self-attention with other mask tokens, its average should be on par with the cross-attention. Yet, the quantitative comparison in Figure 1.(b) shows the magnitude of mask token-to-visible token cross-attention (1.42) in the MAE decoder evaluated over the entire ImageNet validation set far exceeds that of mask token-to-mask token self-attention (0.39).
@@
This initial observation prompts two questions: 1) Is the self-attention mechanism within mask tokens in the decoder truly necessary for effective representation learning? 2) If not, can each patch be independently generated, allowing the reconstruction of only a small subset of masked patches, which in turn, can facilitate faster pretraining without sacrificing downstream performance?
@@
    Cross-attention for decoding. Rather than passing a concatenation of mask and visible tokens to a self-attention decoder, CrossMAE uses mask tokens to query the visible tokens in a cross-attention decoder to reconstruct the masked patches. In this setting, mask tokens incorporate information from the visible tokens but do not interact with other mask tokens, thereby reducing the sequence length for the decoder and cutting down computational costs.
@@
    Partial reconstruction. After the removal of self-attention, given the encoded features from the visible tokens, the decoding of each mask token becomes conditionally independent from one another. This enables the decoding of only a fraction of masked tokens rather than the entire image.
@@
    Inter-block attention. Due to the separation of visible and mask tokens, we are able to use features from different encoder blocks for each decoder block. Empirically, we find solely relying on the last encoder feature map for reconstruction, the design present in MAE, hurts feature learning. We propose a lightweight inter-block attention mechanism that allows the CrossMAE decoder to dynamically leverage a mix of low-level and high-level feature maps from the encoder, improving the learned representation.
Upon training the model, contrary to prior belief, we find that a ViT encoder pretrained with the reconstruction objective can learn a holistic representation of the image regardless of whether self-attention is used in the decoder. Visually, in Fig. 1.(c) and 2, the reconstruction results of CrossMAE are similar to MAE, although CrossMAE can only attend to the visible patches instead of having diffuse attention over all nearby patches. To our surprise, the downstream performance of CrossMAE is on par with MAE, while maintaining a higher efficiency enabled by cross-attention and partial reconstruction. We show that a ViT-B model trained with CrossMAE partial reconstruction achieves a classification accuracy of 83.5% on the ImageNet validation set, which surpasses its full-reconstruction MAE counterpart. In object detection and instance segmentation on COCO, CrossMAE achieves 52.1 AP and 46.3 AP, again surpassing MAE. Finally, we show that with ViT-L, CrossMAE enjoys improved scalability when compared to MAE.
@@ 2.1 Self-Supervised Learning
In self-supervised representation learning (SSL), a model trains on a pretext task where the supervision comes from the input data itself without relying on labeled data. Contrastive learning is a popular approach to SSL that aims to learn representations by contrasting positive and negative samples, such as SimCLR [13], CPC [45], MoCo [30], MoCo v2 [14], BYOL [27], and DINO [10]. Additionally, group-instance contrastive learning works, such as DeepCluster [8], CLD [58] and SwAV [9], integrate clustering into contrastive learning to improve the quality of learned representations.
An alternative method for self-supervised learning is generative modeling, which focuses on acquiring a generative model capable of capturing the underlying data distribution. VAE/GAN [36] merges the strengths of variational autoencoders (VAEs) and generative adversarial networks (GANs) to acquire disentangled representations of data. PixelCNN, PixelVAE, and PixelTransformer [55, 28, 54] generate images pixel by pixel, taking into account the context of previously generated pixels. Masked modeling, a large subclass of generative modeling, is discussed in the following subsection. After the pre-training stage, these generative models can be finetuned for many downstream applications.
@@ 2.2 Masked Modeling
Masked modeling learns representations by reconstructing a masked portion of the input. Pioneering works in natural language processing (NLP) present various such pretraining objectives. BERT [20] and its extensions [42, 35] use a bidirectional transformer and present few-shot learning capabilities from masked language modeling. GPT [47, 48, 6], uses autoregressive, causal masking and demonstrates multi-task, few-shot, and in-context learning capabilities.
@@
Early works in computer vision, such as Stacked Denoising Autoencoders [57] and Context Encoder [46], investigated masked image modeling as a form of denoising or representation learning. Recently, with the widespread use of transformer [21] as a backbone vision architecture, where images are patchified and tokenized as sequences, researchers are interested in how to transfer the success in language sequence modeling to scale vision transformers. BEiT [4], MAE [31], and SimMIM [60] are a few of the early works that explored BERT-style pretraining of vision transformers. Compared to works in NLP, both MAE and SimMIM [31, 60] find that a much higher mask ratio compared to works in NLP is necessary to learn good visual representation. Many recent works further extend masked pretraining to hierarchical architectures [60, 41] and study data the role of data augmentation [11, 22]. Many subsequent works present similar successes of masked pretraining for video [52, 23, 29], language-vision and multi-modal pretraining [2, 40, 24] and for learning both good representations and reconstruction capabilities [59, 38].
@@
However, BERT-style pretraining requires heavy use of self-attention, which makes computational complexity scale as a polynomial of sequence length. PixelTransformer [54] and DiffMAE [59] both use cross-attention for masked image generation and representation learning. Siamese MAE [29] uses an asymmetric masking pattern and decodes frames of a video condition on an earlier frame. In these settings, all masked patches are reconstructed. In this work, we investigate if learning good features necessitates high reconstruction quality and if the entire image needs to be reconstructed to facilitate representation learning. Similar in spirit, PCAE [37] progressively discards redundant mask tokens through its network, leading to a few tokens for reconstruction. In comparison, we minimally modify MAE and start decoding with a random subset of mask tokens.
@@
Figure 3: MAE [31] concatenates all mask tokens with the visible patch features from a ViT encoder and passes them to a decoder with self-attention blocks to reconstruct the original image. Patches that correspond to visible tokens are then dropped, and an L2 loss is applied to the rest of the reconstruction as the pretraining objective. CrossMAE instead uses cross-attention blocks in the decoder to reconstruct only a subset of the masked tokens.
Figure 4: Overview of CrossMAE. (a) The vanilla version of CrossMAE uses the output of the last encoder block as the keys and queries for cross-attention. The first decoder block takes the sum of mask tokens and their corresponding positional embeddings as queries, and subsequent layers use the output of the previous decoder block as queries to reconstruct the masked patches. (b) Unlike the decoder block in [56], the cross-attention decoder block does not contain self-attention, decoupling the generation of different masked patches. (c) CrossMAE’s decoder blocks can leverage low-level features for reconstruction via inter-block attention. It weighs the intermediate feature maps, and the weighted sum of feature maps is used as the key and value for each decoder block.
@@ 2.3 Applications of Cross-Attention
In addition to the prevalent use of self-attention in computer vision, cross-attention has shown to be a cost-effective way to perform pooling from a large set of visible tokens. Intuitively, cross-attention can be seen as a parametric form of pooling, which learnably weighs different features.  [53] replaces mean pooling with cross-attention pooling and finds improvement in ImageNet classification performance. [33] uses cross-attention to efficiently process large volumes of multi-modal data. Cross-attention is also widely used for object detection. [7] utilizes query tokens as placeholders for potential objects in the scene. [17, 16] further extend this concept by introducing additional query tokens to specifically tackle object segmentation in addition to the query tokens for object detection. In this work, we are interested in cross-attention as an efficient method for self-supervised representation learning.
@@ 3 CrossMAE
This section is organized as follows. In Sec. 3.1, we first revisit vanilla Masked Autoencoders. In Sec. 3.2, we propose to use cross-attention instead of self-attention in the decoder for reconstruction. Thanks to a decoding architecture without self-attention, we achieve further efficiency gains by reconstructing only a subset of mask tokens for faster pretraining in Sec. 3.3. Since the use of cross-attention allows different features for different decoder blocks, in Sec. 3.4, we further propose inter-block attention to allow different decoder blocks to focus on different encoder features, which relieves the need to carry all information throughout the neural networks and allows for enhanced feature learning.
@@ 3.1 Preliminaries: Masked Autoencoders
Masked Autoencoders (MAE) [31] pretrain Vision Transformers (ViTs) [21]. Each image input is first patchified, and then a random subset of the patches is selected as the visible patches. As depicted in Fig. 3, the visible patches, concatenated with a learnable class token (<cls>), are subsequently fed into the ViT encoder, which outputs a set of feature latents. The latent vectors, concatenated with the sum of the positional embeddings of the masked patches and the learnable mask token, are passed into the MAE decoder. The decoder blocks share the same architecture as the encoder blocks (i.e., both are transformer blocks with self-attention layers). Note that the number of tokens fed into the decoder is the same length as the original input, and the decoding process assumes that the decoded tokens depend on both visible and masked tokens. Decoder outputs are then passed through a fully connected layer per patch for image reconstruction. After the reconstruction is generated, the loss is applied only to the masked positions, while the reconstructions for visible spatial locations are discarded.
Recall in Sec. 1, to study the properties of MAE, we measure the mean attention value across all attention maps over the ImageNet validation set. We group the attention values by cross-attention and self-attention between visible and masked tokens. We observe that in the decoding process of an MAE, mask tokens attend disproportionately to the class token and the visible tokens (see Figure 1.(b)). This motivates us to make design decisions and conduct experiments specifically to answer the following two questions:
  1. 1.
    Can good representations be learned if masked tokens can attend only to visible tokens?
  2. 2.
    Can we improve pretraining efficiency by reconstructing only part of an image?

### v2

@@ Abstract
In this work, we examine the impact of inter-patch dependencies in the decoder of masked autoencoders (MAE) on representation learning. We decompose the decoding mechanism for masked reconstruction into self-attention between mask tokens and cross-attention between masked and visible tokens. Our findings reveal that MAE reconstructs coherent images from visible patches not through interactions between patches in the decoder but by learning a global representation within the encoder. This discovery leads us to propose a simple visual pretraining framework: cross-attention masked autoencoders (CrossMAE). This framework employs only cross-attention in the decoder to independently read out reconstructions for a small subset of masked patches from encoder outputs. This approach achieves comparable or superior performance to traditional MAE across models ranging from ViT-S to ViT-H and significantly reduces computational requirements. By its design, CrossMAE challenges the necessity of interaction between mask tokens for effective masked pretraining. Code and models are publicly available: https://crossmae.github.io/.
footnotetext: Equal contribution. Equal advising.
@@ 1 Introduction
Masked image modeling (Pathak et al., 2016; He et al., 2022; Xie et al., 2022; Bao et al., 2022) has emerged as a pivotal unsupervised learning technique in computer vision. One such recent work following this paradigm is masked autoencoders (MAE): given only a small, random subset of visible image patches, the model is tasked to reconstruct the missing pixels. By operating mostly on this small subset of visible tokens, MAE can efficiently pre-train high-capacity models on large-scale vision datasets, demonstrating impressive results on a wide array of downstream tasks (Kirillov et al., 2023; Li et al., 2022b; Radosavovic et al., 2023).
The MAE framework employs self-attention across the entire model for self-supervised reconstruction tasks. In this setup, both masked and visible tokens engage in self-attention, not just with each other but also with themselves, aiming to generate a holistic and context-aware representation. However, the masked tokens inherently lack information. Intuitively, facilitating information exchange among adjacent masked tokens should enable the model to synthesize a more coherent image, thereby accomplishing the task of masked reconstruction and improving representation learning. A question arises, though: Is this truly the case?
@@
We decompose the decoding process of each mask token into two parallel components: self-attention with other mask tokens, as well as cross-attention to the encoded visible tokens. If MAE relies on the self-attention with other mask tokens, its average should be on par with the cross-attention. Yet, the quantitative comparison in Figure 1.(b) shows the magnitude of mask token-to-visible token cross-attention (1.42) in the MAE decoder evaluated over the entire ImageNet validation set far exceeds that of mask token-to-mask token self-attention (0.39).
Figure 1: Method Overview. (A) Masked autoencoder (MAE) starts by masking random patches of the input image. (B) To reconstruct a mask token (marked by the blue star), MAE attends to both the masked tokens (B.Left) and the visible tokens (B.Right). A quantitative comparison over the ImageNet validation set shows that the masked tokens in MAE disproportionally attend to the visible tokens (1.42 vs 0.39), questioning the necessity of attention within mask tokens. (C) We propose CrossMAE, the masked patches are reconstructed from only the cross attention between the masked tokens and the visible tokens. Surprisingly, CrossMAE attains the same or better performance than MAE on ImageNet classification and COCO instance segmentation.
Figure 2: Example reconstructions of ImageNet validation images. For each set of 5 images, from left to right, are the original image, masked image with a mask ratio of 75%, MAE He et al. (2022), CrossMAE (trained to reconstruct 25% of image tokens, or 1/3 of the mask tokens), and CrossMAE (trained to reconstruct all masked tokens). Since CrossMAE does not reconstruct them, all model outputs have the visible patches overlaid. Intriguingly, CrossMAE, when trained for partial reconstruction, can decode all mask tokens in one forward pass (shown above), indicating that the encoder rather than the decoder effectively captures global image information in its output tokens. Its comparable reconstruction quality to full-image-trained models suggests that full-image reconstruction might not be essential for effective representation learning.
@@
This initial observation prompts two questions: 1) Is the self-attention mechanism among mask tokens in the decoder necessary for effective representation learning? 2) If not, can each patch be independently read out from the encoder output, allowing the reconstruction of only a small subset of masked patches, which in turn, accelerates the pretraining without performance degradation?
@@
    Cross-attention for decoding. Rather than passing a concatenation of mask and visible tokens to a self-attention decoder, CrossMAE uses mask tokens as queries to read out the masked reconstructions from the visible tokens in a cross-attention decoder. In this setting, mask tokens incorporate information from the visible tokens but do not interact with other mask tokens, thereby reducing the sequence length for the decoder and cutting down computational costs.
@@
    Independent partial reconstruction. With self-attention removed, the decoding of each mask token, based on the encoded features from visible tokens, becomes conditionally independent. This enables the decoding of only a fraction of masked tokens rather than the entire image.
@@
    Inter-block attention. Due to the separation of visible and mask tokens, we can use features from different encoder blocks for each decoder block. Empirically, we find solely relying on the last encoder feature map for reconstruction, the design present in MAE, hurts feature learning. We propose a lightweight inter-block attention mechanism that allows the CrossMAE decoder to leverage a mix of low-level and high-level feature maps from the encoder, improving the learned representation.
The analysis performed on CrossMAE led to a novel way to understand MAE. Even though the patches to be reconstructed are independently decoded, our findings demonstrate that coherent reconstruction for each masked patch can be independently read out from the encoder output, without any interactions among masked tokens in the decoder for consistency (Figure 2). Furthermore, the downstream performance of the model remains robust even without these interactions (Figure 1.(c), Tables 1 and 2). Both pieces of evidence confirm that the encoder’s output features already encapsulate the necessary global context for image reconstruction, while the decoder simply performs a readout from the encoder output to reconstruct the pixels at the location of each patch.
To sum up, our main contributions are the following:
  1. 1.
    We present a novel understanding of MAE. Our findings show that MAE reconstructs coherent images from visible patches not through interactions between patches to be reconstructed in the decoder but by learning a global representation within the encoder. This is evidenced by the model’s ability to generate coherent images and maintain robust downstream performance without such interactions, indicating the encoder effectively captures global image information.
  2. 2.
    We advocate replacing self-attention layers with a simple cross-attention readout function. Given our discovery that the encoder in MAE already captures a comprehensive global representation, we propose replacing self-attention layers in the decoder with a more efficient information readout function. Specifically, we suggest utilizing cross-attention to aggregate the output tokens of the encoder into each input token within the decoder layers independently, thereby eliminating the need for token-to-token communication within the decoder.
  3. 3.
    CrossMAE achieves comparable or superior performance with reduced computational costs in image classification and instance segmentation compared to MAE on vision transformer models ranging from ViT-S to ViT-H. Code is available here.
@@ 2.1 Self-Supervised Learning
In self-supervised representation learning, a model trains on a pretext task where the supervision comes from the input data itself without labels. Contrastive learning methods learn representations by contrasting positive and negative samples, such as SimCLR (Chen et al., 2020b), CPC (Oord et al., 2018), MoCo (He et al., 2020; Chen et al., 2020c; 2021a), CLD (Wang et al., 2021) and SwAV (Caron et al., 2020). Additionally, in BYOL (Grill et al., 2020), iBOT (Zhou et al., 2021), DINO (Caron et al., 2021), DINOv2 (Oquab et al., 2023), and MaskAlign (Xue et al., 2023) make a student model to imitate a teacher model without negative pairs.
Generative modeling, focusing on acquiring a generative model capable of capturing the underlying data distribution, is an alternative method for self-supervised learning. VAE/GAN (Larsen et al., 2016) merges the strengths of variational autoencoders and generative adversarial networks to acquire disentangled representations of data. PixelCNN, PixelVAE, and PixelTransformer (Van den Oord et al., 2016; Gulrajani et al., 2016; Tulsiani & Gupta, 2021) generate images pixel by pixel, taking into account the context of previously generated pixels. Masked modeling, a large subclass of generative modeling, is discussed in the following subsection. After the pre-training stage, these generative models can be finetuned for many downstream applications.
@@ 2.2 Masked Modeling
Masked modeling learns representations by reconstructing a masked portion of the input. Pioneering works in natural language processing (NLP) present various such pretraining objectives. BERT (Devlin et al., 2019) and its extensions (Liu et al., 2019; Lan et al., 2020) use a bidirectional transformer and present few-shot learning capabilities from masked language modeling. GPT (Radford et al., 2018; 2019; Brown et al., 2020), uses autoregressive, causal masking and demonstrates multi-task, few-shot, and in-context learning capabilities.
@@
Early works in computer vision, such as Stacked Denoising Autoencoders (Vincent et al., 2010) and Context Encoder (Pathak et al., 2016), investigated masked image modeling as a form of denoising or representation learning. Recently, with the widespread use of transformer (Dosovitskiy et al., 2020) as a backbone vision architecture, where images are patchified and tokenized as sequences, researchers are interested in how to transfer the success in language sequence modeling to scale vision transformers. BEiT (Bao et al., 2021), MAE (He et al., 2022), and SimMIM (Xie et al., 2022) are a few of the early works that explored BERT-style pretraining of vision transformers. Compared to works in NLP, both MAE and SimMIM (He et al., 2022; Xie et al., 2022) find that a much higher mask ratio compared to works in NLP is necessary to learn good visual representation. Many recent works further extend masked pretraining to hierarchical architectures (Xie et al., 2022; Liu et al., 2022) and study data the role of data augmentation (Chen et al., 2023; Fang et al., 2023). Many subsequent works present similar successes of masked pretraining for video (Tong et al., 2022; Wang et al., 2023; Feichtenhofer et al., 2022; Gupta et al., 2023), language-vision and multi-modal pretraining (Bachmann et al., 2022; Li et al., 2023b; Geng et al., 2022) and for learning both good representations and reconstruction capabilities (Wei et al., 2023; Li et al., 2022a).
@@
However, BERT-style pretraining requires heavy use of self-attention, which makes computational complexity scale as a polynomial of sequence length. PixelTransformer (Tulsiani & Gupta, 2021) and DiffMAE (Wei et al., 2023) both use cross-attention for masked image generation and representation learning. Siamese MAE (Gupta et al., 2023) uses an asymmetric masking pattern and decodes frames of a video condition on an earlier frame. In these settings, all masked patches are reconstructed. In this work, we investigate if learning good features necessitates high reconstruction quality and if the entire image needs to be reconstructed to facilitate representation learning. PCAE (Li et al., 2023a) progressively discards redundant mask tokens through its network, leading to a few tokens for reconstruction. VideoMAEv2 (Wang et al., 2023) concatenates randomly sampled masked tokens with visible tokens and uses self-attention to reconstruct the masked patches. In comparison, we minimally modify MAE with a cross-attention-only decoder and masked tokens are decoded in a conditional independent way.
@@
Figure 3: MAE He et al. (2022) concatenates all mask tokens with the visible patch features from a ViT encoder and passes them to a decoder with self-attention blocks to reconstruct the original image. Patches that correspond to visible tokens are then dropped, and an L2 loss is applied to the rest of the reconstruction as the pretraining objective. CrossMAE instead uses cross-attention blocks in the decoder to reconstruct only a subset of the masked tokens.
@@ 2.3 Applications of Cross-Attention
In addition to the prevalent use of self-attention in computer vision, cross-attention has shown to be a cost-effective way to perform pooling from a large set of visible tokens. Intuitively, cross-attention can be seen as a parametric form of pooling, which learnably weighs different features. Touvron et al. (2021) replace mean pooling with cross-attention pooling and find improvement in ImageNet classification performance. Jaegle et al. (2021) uses cross-attention to efficiently process large volumes of multi-modal data. Cross-attention is also widely used for object detection. Carion et al. (2020) utilizes query tokens as placeholders for potential objects in the scene. Cheng et al. (2022; 2021) further extend this concept by introducing additional query tokens to specifically tackle object segmentation in addition to the query tokens for object detection. Distinct from thes prior works, we are interested the role of cross-anttention for representation learning in a self-supervised manner.
@@ 3 CrossMAE
We start with an overview of vanilla masked autoencoders in Section 3.1. Next, in Section 3.2, we introduce the use of cross-attention in place of self-attention in the decoder for testing the necessity of interaction between mask tokens for representation learning. In Section 3.3, we discuss how eliminating self-attention in the decoding process enables us to reconstruct only a subset of masked tokens, leading to faster pretraining. Finally, Section 3.4 presents our inter-block attention mechanism, which allows decoder blocks to leverage varied encoder features.
Figure 4: Overview of CrossMAE. (a) The vanilla version of CrossMAE uses the output of the last encoder block as the keys and queries for cross-attention. The first decoder block takes the sum of mask tokens and their corresponding positional embeddings as queries, and subsequent layers use the output of the previous decoder block as queries to reconstruct the masked patches. (b) Unlike the decoder block in Vaswani et al. (2017), the cross-attention decoder block does not contain self-attention, decoupling the generation of different masked patches. (c) CrossMAE’s decoder blocks can leverage low-level features for reconstruction via inter-block attention. It weighs the intermediate feature maps, and the weighted sum of feature maps is used as the key and value for each decoder block.
@@ 3.1 Preliminaries: Masked Autoencoders
Masked Autoencoders (MAE) (He et al., 2022) pretrain Vision Transformers (ViTs) (Dosovitskiy et al., 2020). Each image input is first patchified, and then a random subset of the patches is selected as the visible patches. As depicted in Figure 3, the visible patches, concatenated with a learnable class token [CLS], are subsequently fed into the ViT encoder, which outputs a set of feature latents. The latent vectors, concatenated with the sum of the positional embeddings of the masked patches and the learnable mask token, are passed into the MAE decoder. The decoder blocks share the same architecture as the encoder blocks (i.e., both are transformer blocks with self-attention layers). Note that the number of tokens fed into the decoder is the same length as the original input, and the decoding process assumes that the decoded tokens depend on both visible and masked tokens. Decoder outputs pass through a fully connected layer per patch for image reconstruction. After the reconstruction is generated, the loss is applied only to the masked positions, while the reconstructions for visible spatial locations are discarded.
Recall in Sec. 1 we measure the mean attention value across all attention maps over the ImageNet validation set to study the properties of MAE. We grouped the attention values by cross-attention and self-attention between visible and masked tokens. We observed that in the decoding process of an MAE, mask tokens attend disproportionately to the class token and the visible tokens (see Figure 1.(b)). This motivates us to make design decisions and conduct experiments specifically to answer the following question: Can we simplify the decoding process by eliminating self-attention among masked tokens without compromising the model’s ability to generate coherent images and perform well on downstream tasks?